OpenStudy (anonymous):

Medal & fan. Perform the indicated operation. Be sure the answer is reduced. (equation below)

OpenStudy (anonymous):

\[\frac{ k+5 }{ k^2+3k-10 } \div \frac{ 7k+14 }{ 4-k^2 }\]

OpenStudy (anonymous):

@taramgrant0543664

OpenStudy (taramgrant0543664):

So since its divided by a division you can flip the second portion so it is a multiplication so it would be x(4-k^2)/(7k+14)

OpenStudy (taramgrant0543664):

Our next step is to make k^2+3k-10 into something that looks more like ( )( ) obviously with numbers in it it but we need to determine that

OpenStudy (taramgrant0543664):

If factored properly you should get -2 and 5 that solve for k^2+3k-10 so k^2+3k-10 = (k-2)(k+5)